# Boosting Ensembles for Statistics of Tails at Conditionally Optimal Advance Split Times

### v1

@@ 2 Sampling and estimation methodology
    estimate conditional tail distributions for each event and each AST separately;
@@
    combine the conditional tails into an unconditional (“climatological”) tail, using the estimators specified below, for a range of ASTs, and select the optimal one based on the skill of the corresponding tail estimate in reproducing the tail of a “long DNS”.
    We then display the results of applying this procedure is then repeated across a range of target locations in the model flow domain,
Figure 1: Schematic summarizing the ensemble boosting and tail estimation procedure, using a simple Langevin dynamics with a potential that is quadratic for $x\in(-0.25,0.25)$—the blue-shaded region—and logarithmic outside this range. The system exhibits intermittent, transient extremes (a.i) and a power law tail $\mathbb{P}\{X>x\}\sim x^{-3.1}$ (a.ii). We set a threshold for severity (horizontal black dashed line) at roughly the minimum probability estimable from the relatively short (duration 1600) timeseries (see the black empirical PDF in a.ii and the black empirical CCDFs in (b,c,d).iii, as compared with the true PDF and CCDF in gray). We then identify the peaks over the threshold (vertical black dashed lines in a.i), and perturb the simulation in advance of these peaks. Three choices of advance split time (AST) are shown in rows b,c,d, marked by vertical red lines, each resulting in “boosted” peak distributions (red curves in b,c,d).(i,ii), described by complementary CDFs (CCDFs) shown in light red in (b,c,d).(iii). Combining these conditional CCDFs together using the “MoCTail” estimator introduced later in Eq. (16) gives the dark red dashed line, which is meant to approximate the ground truth (gray line) better than the short DNS alone can do. The intermediate AST (c) is best among the three for this task, and our goal is to formulate and characterize this optimal AST more generally.
@@ 2.1 Generating the dataset of boosted ensembles
We run a direct numerical simulation (“short DNS”) $\{\mathbf{x}(t):0\leq t\leq T_{\text{short}}\}$, long enough to generate some extremes but not enough to estimate probabilities smaller than $1/(\epsilon^{2}T_{\text{short}})=100/T_{\text{short}}$ for a relative error tolerance of $\epsilon=0.1$. The premise of RES, and ensemble boosting, is that the extremes it does generate might have been even worse, perhaps just a butterfly flap away from the more intense extremes one would see with a “long DNS” of duration $T_{\text{long}}\gg T_{\text{short}}$. We generate the long DNS as well to serve as a ground-truth for validation. Following the ensemble boosting methodology laid out in Gessner et al., (2021); Gessner, (2022); Fischer et al., (2023) and Noyelle, (2024), we first identify a threshold $\mu$ with exceedance probability $q(\mu)$ that is moderate enough to estimate precisely with the short DNS. In other words, $\mu$ is the $[1-q(\mu)]$th quantile, or “$q(\mu)$th complementary quantile”. Equivalently, $q(\mu)$ is the complementary cumulative density function (CCDF) of the random variable $R$, evaluated at $\mu$. In line with the peaks-over-threshold procedure (Coles,, 2001), we take cluster maxima of exceedances above $\mu$ as the “ancestral” extreme events. Concretely, a cluster maximum is a state from the DNS, $\mathbf{x}^{*}=\mathbf{x}(t^{*})$, such that
@@
and for a sequence of increasing ASTs $\{A_{j}:j=1,\ldots,J\}$ bounded between 0 and $A_{\text{max}}$, launch an ensemble of descendants $\{\mathbf{x}^{*}_{n,j,m}:m=1,\ldots,M_{n,j}\}$ generated by applying $M_{n,j}$ different perturbations to the DNS at time $t_{n}^{*}-A_{j}$, and running each simulation to time $t_{n}^{*}+B$. Note that $M_{n,j}$ could in principle vary between ancestors $n$ and lead times $j$, which is not needed for our exhaustive sweeps in this paper, but certainly would be needed in an “online” rare event sampling procedure that iteratively homes in on a subset of the most extreme-ogenic ancestors $\{n\}$ and ASTs $\{j\}$ to draw more samples from.
[…]
@@
The other common case is where $\mathbf{x}(t)$ is a stochastic process, e.g., an Ito diffusion forced by white noise, as we used in Finkel and O’Gorman, (2024) as well as the schematic in Fig. 1. In that case, $\omega$ is a white noise process sampled at discrete times, whose dimensionality scales with the number of timesteps. In the QG experiments, we adhere to impulsive perturbations for three reasons: it introduces fewer arbitrary parameters, it is less disruptive to the system’s intrinsic dynamics, and it keeps the dimensionality of the random space low. If, as we conjecture, even low-dimensional butterfly flaps are sufficient to excite the more extreme fluctuations, it would make deterministic search methods—which should always be preferred over Monte Carlo—more viable.
Following the perturbation, the descendant drifts away from the parent and achieves its own severity $R^{*}$ when its intensity function $R$ peaks at some time $t_{n,j,m}^{*}$ possibly different from its ancestor’s peak time $t_{n}^{*}$:
@@
[…]
@@ 2.2 Estimating conditional and climatological probabilities from boosted ensembles
Assume now there is a measure $\mathbb{P}$ on $\Omega$ with associated density function $p(\omega)$, which might for example place higher weight on smaller kicks. Each ensemble of descendants at each lead time gives rise to its own conditional severity distribution (as opposed to “climatological,” due to its association with the $n$th ancestor’s particular initial condition):
@@
which can be estimated from the samples $\{R_{n,j,m}^{*}:m=1,\ldots,M_{n,j}\}$. Whereas Monte Carlo is the typical strategy in rare event sampling (Finkel and O’Gorman,, 2024; Bloin-Wibe et al.,, 2025), the deliberately low-dimensional perturbations that we employ here enable numerical quadrature instead. Based on the samples, we fit some parametric model $\widehat{R}_{n,j}^{*}(\omega;\theta)$ with parameters $\theta$, for example polynomial coefficients, kernel weights, or neural network weights. Then the integral over $\Omega$ can be estimated, either analytically (if $p$ and $\widehat{R}^{*}$ take simple enough forms) or numerically by densely filling $\Omega$ with a grid of points, evaluating $\widehat{R}^{*}$ and $p$ at each point, and taking their inner product. The result is an estimate $\widehat{Q}_{n,j}(r)$ which is found by replacing $R_{n,j}^{*}(\omega)$ with $\widehat{R}_{n,j}^{*}(\omega)$ in Eq. (9).
The tail part of the CCDF above $\mu$ is given by
$Q_{n,j}(r;\mu)=\mathbb{P}\{R_{n,j}^{*}>r|R_{n,j}^{*}>\mu\}=\frac{Q_{n,j}(r)}{Q_{n,j}(\mu)},$ (10)
and can be estimated it by putting hats $\widehat{(\cdot)}$ on every $Q$. However, this risks dividing by zero, because the fitted function $\widehat{Q}_{n,j}$ may imply zero probability of exceeding the threshold, particularly at long ASTs when descendants have enough time to decorrelate totally with their ancestor. To prevent this, we implement a continuous version of the “accept-reject” from TEAMS procedure, replacing the PDF $p(\omega)$ over all regions of $\Omega$ where $\widehat{R}^{*}(\omega)<\mu$ (which spawns “rejected descendants”) with the Dirac delta measure $\delta_{0}(\omega)$ (which, by definition spawns the ancestor):
$\widehat{Q}_{n,j}(r;\mu)$ :=$:=\begin{cases}\widehat{Q}_{n,j}(r)&\text{if }\widehat{Q}_{n,j}(\mu)>0\\ \mathbb{I}\{R_{n}^{*}>r\}&\text{ otherwise}\end{cases}$ (11)
$=\widehat{Q}_{n,j}(r)+\mathbb{I}\{R_{n}^{*}>r\}\big{[}1-\widehat{Q}_{n,j}(\mu)\big{]}$ (12)
($\widehat{Q}_{n,j}(r)=0$ when $\widehat{Q}_{n,j}(\mu)=0$ since $Q_{n,j}$ is decreasing, hence the two terms in the last expression correspond to the two cases). Another heuristic way to justify this expression is to stipulate that we care about approximating only the extreme part of the boosting distribution, i.e., those $\omega$ near enough to 0 that $R^{*}(\omega)>\mu$, hence $\widehat{Q}(\mu)$ is close to 1, allowing for the Taylor expansion
$Q_{n,j}(r;\mu)$ $=\frac{Q_{n,j}(r)}{Q_{n,j}(\mu)}=\frac{Q_{n,j}(r)}{1-[1-Q_{n,j}(\mu)]}\approx Q_{n,j}(r)+Q_{n,j}(r)[1-Q_{n,j}(\mu)]$ (13)
$\approx\widehat{Q}_{n,j}(r)+\mathbb{I}\{R_{n}^{*}>r\}\big{[}1-\widehat{Q}_{n,j}(\mu)\big{]}$ (14)
$=:\widehat{Q}_{n,j}(r;\mu)$ (15)
@@
We then estimate the unconditional (climatological) CCDF as a uniform mixture over ancestors, selecting one representative AST $A_{j_{n}}$ from each ancestor $n$ to best represent its alternate realities according to some selection rule (different rules will be evaluated thoroughly for the QG system in Sect. 6).
$\widehat{Q}^{M}(r;\mu)=\frac{1}{N_{\text{short}}}\sum_{n=1}^{N_{\text{short}}}\widehat{Q}_{n,j_{n}}(r;\mu).$ (16)
We call this the “MoCTail” estimator, for “Mixture of Conditional Tails.”
@@
The recent works Noyelle, (2024) and Bloin-Wibe et al., (2025) formulate a different estimator, which makes for an interesting comparison. Rather than summing $N_{\text{short}}$ tail CCDFs, each approximating a ratio of the form (10), they construct a single ratio by summing $N_{\text{short}}$ numerators and $N_{\text{short}}$ denominators. Translated into our own notation, this becomes
$\widehat{Q}^{P}(r;\mu)=\frac{\sum_{n=1}^{N_{\text{short}}}\widehat{Q}_{n,j_{n}}(r)}{\sum_{n=1}^{N_{\text{short}}}\widehat{Q}_{n,j_{n}}(\mu)}.$ (17)
We call this the “PoPTail” estimator, for “Pool of Perturbed Tails.”
One could argue for either estimator based on the validity of its underlying assumptions. Bloin-Wibe et al., (2025) develop the PoPTail estimator  (17) by arguing that the (numerator, denominator) estimate conditional probabilities $\mathbb{P}\{R^{*}>(r,\mu)|\text{AC}_{t}^{\epsilon}\}$, where $\text{AC}_{t}^{\epsilon}$ is the set of states $\epsilon$-close to initial conditions that will lead to exceeding $\mu$; however, it assumes that the DNS only passed through $\text{AC}_{t}^{\epsilon}$ on its way to an actual threshold-crossing event. It neglects the possibility of “near misses”: times from the DNS run that would have reached $\mu$ but for an $\epsilon$-perturbation, and missed the chance to become ancestors. We suspect it is more harder to justify either estimator on airtight mathematical grounds than Bloin-Wibe et al., (2025) suggest, and here adopt a more openly empirical perspective in testing the skill of both. We do this based on the $\chi^{2}$-divergence against the “ground truth” $Q$ as estimated by a long DNS: with a sequence of thresholds $\mu=r_{0}<r_{1}<r_{2}<\ldots<r_{K-1}<r_{K}=\infty$, and defining the probability mass function $\Delta Q_{k}=Q_{k}-Q_{k+1}$ as the probability contained in the $k$th bin (note that $Q_{K}=0$ and so $\Delta Q_{K-1}=Q_{K-1}$), the $\chi^{2}$-divergence of either estimator $\widehat{Q}\in\{\widehat{Q}^{M},\widehat{Q}^{p}\}$ is defined as
$\chi^{2}(\Delta\widehat{Q}\lVert\Delta Q)=\sum_{k=0}^{K-1}\frac{(\Delta Q_{k}-\Delta\widehat{Q}_{k})^{2}}{\Delta Q_{k}}$ (18)

### v2

@@ 2 Sampling and estimation methodology
    estimate tail distributions, conditional on the event and the AST;
@@
    combine the conditional tails into an unconditional (“climatological”) tail, using the estimators specified below, for a range of ASTs, and select the optimal AST based on the skill of the corresponding tail estimate in reproducing the tail of a “long DNS”.
We then display the results of applying this procedure to a range of target locations in the model flow domain.
Figure 1: Schematic summarizing the ensemble boosting and tail estimation procedure, using a simple Langevin dynamics with a potential that is quadratic for $x\in(-0.25,0.25)$—the blue-shaded region—and logarithmic outside this range. Appendix A specifies the system completely. The position variable $X(t)$ exhibits intermittent, transient extremes (a.i) and power law tails $\mathbb{P}\{|X|>|x|\}\sim|x|^{-3.1}$ (a.ii). We set a threshold for severity (horizontal black dashed line) at roughly the minimum probability estimable from a relatively short (duration 1600) timeseries (see the black empirical PDF in a.ii and the black empirical CCDFs in (b,c,d).iii, as compared with the true PDF and CCDF in gray). We then identify the peaks over the threshold (marked by vertical black dashed lines in a.i), and perturb the simulation in advance of these peaks. Three choices of advance split time (AST) are shown in rows b,c,d, marked by vertical red lines, each resulting in “boosted” peak ensembles, shown as red curves in (b,c,d).(i,ii) and summarized by complementary CDFs (CCDFs) shown in light red in (b,c,d).(iii). Combining these conditional CCDFs together using the “MoCTail” estimator introduced later in Eq. (18) gives the dark red dashed line, which is meant to approximate the ground truth (gray line) better than the short DNS alone can do, including by going to higher values of $x$. The intermediate AST (c) is best among the three for this task, and our goal is to formulate and characterize this optimal AST more generally.
@@ 2.1 Generating the dataset of boosted ensembles
There are many design choices in ensemble boosting (Gessner et al., 2021): how to select extreme events to boost, how many boosts to generate, when to launch them, etc. This subsection details the choices used here.
We run a direct numerical simulation (“short DNS”) $\{\mathbf{x}(t):0\leq t\leq T_{\text{short}}\}$, long enough to generate some extremes but not enough to estimate probabilities smaller than $1/(\epsilon^{2}T_{\text{short}})=100/T_{\text{short}}$ for a relative error tolerance of $\epsilon=0.1$. The premise of RES, and ensemble boosting, is that the extremes it does generate might have been even worse, perhaps just a butterfly flap away from the more intense extremes one would see with a “long DNS” of duration $T_{\text{long}}\gg T_{\text{short}}$. We generate such a long DNS as well to serve as a ground-truth for validation. Following the ensemble boosting methodology laid out in Gessner et al. (2021); Gessner (2022); Fischer et al. (2023) and Noyelle (2024), we first identify a threshold $\mu$ with exceedance probability $q(\mu)$ that is moderate enough to estimate precisely with the short DNS. In other words, $\mu$ is the $1-q(\mu)]$th quantile, or $q(\mu)$th complementary quantile”. Equivalently, $q(\mu)$ is the complementary cumulative density function (CCDF) of the random variable $R$, evaluated at $\mu$. In line with the peaks-over-threshold procedure (Coles, 2001), we take cluster maxima of exceedances above $\mu$ as the “ancestral” extreme events. Concretely, a cluster maximum is a state from the DNS, $\mathbf{x}^{*}=\mathbf{x}(t^{*})$, such that
@@
and for a sequence of increasing ASTs $\{A_{j}:j=1,\ldots,J\}$ bounded between 0 and $A_{\text{max}}$, launch an ensemble of descendants $\{\mathbf{x}^{*}_{n,j,m}:m=1,\ldots,M_{n,j}\}$ by applying $M_{n,j}$ different perturbations to the DNS at time $t_{n}^{*}-A_{j}$, and running each simulation to time $t_{n}^{*}+B$. Note that $M_{n,j}$ could in principle vary between ancestors $n$ and lead times $j$, which is not needed for our exhaustive sweeps in this paper, but certainly would be needed in an “online” rare event sampling procedure that iteratively homes in on a subset of the most extreme-ogenic ancestors $\{n\}$ and ASTs $\{j\}$ to draw more samples from.
A bit more notation helps clarify how the perturbing is done, abstractly at first and concretely in Sect. 3 when we specialize to the QG system. For each $(n,j,m)$, we draw a random sample $\omega_{n,j,m}$ from some sample space $\Omega$. Denoting $\Phi^{\Delta t}:\mathbb{R}^{d}\times\Omega\to\mathbb{R}^{d}$ be the flow map that integrates the perturbed dynamics forward by a time interval $\Delta t$, the $(n,j,m)$th descendant’s trajectory through state space $\mathbb{R}^{d}$ can be written
@@
The other common case is where $\mathbf{x}(t)$ is a stochastic process, e.g., an Ito diffusion forced by white noise, as we used in Finkel and O’Gorman (2024) as well as the schematic in Fig. 1. In that case, $\omega$ is a white noise process sampled at discrete times, whose dimensionality scales with the number of timesteps. In the QG experiments, we adhere to impulsive perturbations for three reasons: it introduces fewer arbitrary parameters, it is less disruptive to the system’s intrinsic dynamics, and it keeps the dimensionality of the random space low. If, as we conjecture, even low-dimensional butterfly flaps are sufficient to excite the more extreme fluctuations, it would make deterministic search methods—which should always be preferred over Monte Carlo—more viable.
Following the perturbation, the descendant drifts away from the parent and achieves its own severity $R^{*}$ (peak of its intensity function $R$) at some time $t_{n,j,m}^{*}$ possibly different from its ancestor’s peak time $t_{n}^{*}$:
@@
where the latter notation emphasizes dependence on $\omega$, while recognizing that each $(n,j)$ induces a different severity function $R^{*}$ because perturbations may be felt differently depending on the initial condition.
@@ 2.2 Estimating conditional and climatological probabilities from boosted ensembles
Assume now there is a probability measure $\mathbb{P}^{\Omega}$ on $\Omega$ with associated density function $p^{\Omega}(\omega)$, which might for example place higher weight on smaller kicks. The $\Omega$ superscript will generally relate to statistics over this conditional probability measure, to distinguish it from long-term climatological statistics. A major aim of this paper is to show how they relate to each other. Each ensemble of descendants at each lead time gives rise to its own conditional severity distribution:
$Q_{n,j}^{\Omega}(r)=\mathbb{P}^{\Omega}\{R_{n,j}^{*}>r\}=\int_{\Omega}\mathbb{I}\{R_{n,j}^{*}(\omega)>r\}p^{\Omega}(\omega)\,d\omega,$ (5)
which can be estimated from the samples $\{R_{n,j,m}^{*}:m=1,\ldots,M_{n,j}\}$. Here conditional means starting with a perturbation of the $n$th ancestor’s particular initial condition at time $t_{n}^{*}-A_{j}$ and running forward until time $t_{n}^{*}+B$. By contrast, we refer to the climatological severity distribution as that resulting from a long DNS. Whereas Monte Carlo is the typical strategy in rare event sampling due to an imposed, high-dimensional perturbation space meant to represent extrinsic uncertainty (e.g., wind or waves buffeting an engineered structure; Au and Beck, 2001; Mohamad and Sapsis, 2018), in our setting the perturbation space is an arbitrary design choice aiming at an indirect goal (climate estimation), and nothing stops us here from deliberately choosing low-dimensional perturbations instead of high-dimensional ones as in Ragone et al. (2018); Bloin-Wibe et al. (2025). This enables numerical quadrature instead of Monte Carlo, and saves on cost by allowing sample re-use across different input distributions. Determining whether this works is one central question this paper aims to answer.
Based on the samples drawn from $\Omega$, we fit a regression model $\widehat{R}_{n,j}^{*}(\omega;\theta)$ with parameters $\theta$, in our case coefficients for linear and quadratic polynomials. In general $\widehat{R}_{n,j}$ could be a more elaborate parametric model, e.g., a Gaussian process or neural network with learned weights $\theta$, as often used in modern uncertainty quantification (Kabir et al., 2018; Sapsis, 2020; Pickering et al., 2022). Then the integral over $\Omega$ can be estimated, either analytically (if $p$ and $\widehat{R}^{*}$ take simple enough forms) or numerically by densely filling $\Omega$ with a grid of points, evaluating $\widehat{R}^{*}$ and $p$ at each point, and taking their inner product. The result is an estimate $\widehat{Q}^{\Omega}_{n,j}(r)$ for the conditional tail CCDF
$Q^{\Omega}_{n,j}(r;\mu)=\mathbb{P}^{\Omega}\{R_{n,j}^{*}>r|R_{n,j}^{*}>\mu\}=\frac{Q_{n,j}^{\Omega}(r)}{Q_{n,j}^{\Omega}(\mu)},$ (6)
and can be estimated it by putting hats $\widehat{(\cdot)}$ on every $Q$. However, this risks dividing by zero, because the fitted function $\widehat{Q}_{n,j}$ may imply zero probability of exceeding the threshold, particularly at long ASTs when descendants have enough time to decorrelate totally with their ancestor. This loss of ancestral “wisdom” is a more fundamental problem than the numerical issue of zero denominator, and we address it by implementing a continuous version of the “accept-reject” step of the TEAMS procedure in Finkel and O’Gorman (2024). Wherever the descendant severity $\widehat{R}_{n,j}^{*}(\omega)$ falls below $\mu$, we replace it with the ancestor severity, denoted $R_{n}^{*}$ (with no second subscript):
$\widehat{Q}_{n,j}^{\Omega}(r;\mu)$ $:=\int_{\Omega}\left\{\begin{array}[]{ll}\mathbb{I}\{\widehat{R}_{n,j}^{*}(\omega)>r\}&\text{if }\widehat{R}_{n,j}^{*}(\omega)>\mu\\ \mathbb{I}\{R_{n}^{*}>r\}&\text{ otherwise}\end{array}\right\}p^{\Omega}(\omega)\,d\omega$ (9)
$=\int_{\{\widehat{R}_{n,j}^{*}(\omega)>\mu\}}\mathbb{I}\{\widehat{R}_{n,j}^{*}(\omega)>r\}p^{\Omega}(\omega)\,d\omega+\int_{\{\widehat{R}_{n,j}^{*}(\omega)\leq\mu\}}\mathbb{I}\{R_{n}^{*}>r\}p^{\Omega}(\omega)\,d\omega$ (10)
$=\int_{\Omega}\mathbb{I}\{\widehat{R}_{n,j}^{*}(\omega)>r\}p^{\Omega}(\omega)\,d\omega+\mathbb{I}\{R_{n}^{*}>r\}\int_{\{\widehat{R}_{n,j}^{*}(\omega)\leq\mu\}}p^{\Omega}(\omega)\,d\omega$ (11)
$=\widehat{Q}_{n,j}^{\Omega}(r)+\mathbb{I}\{R_{n}^{*}>r\}\big[1-\widehat{Q}_{n,j}^{\Omega}(\mu)\big]$ (12)
@@
$Q_{n,j}^{\Omega}(r;\mu)$ $=\frac{Q_{n,j}^{\Omega}(r)}{Q_{n,j}^{\Omega}(\mu)}$ (13)
$=\frac{Q_{n,j}^{\Omega}(r)}{1-[1-Q_{n,j}^{\Omega}(\mu)]}$ (14)
$\approx Q_{n,j}^{\Omega}(r)+[1-Q_{n,j}^{\Omega}(\mu)]Q_{n,j}^{\Omega}(r)$ (15)
$\approx\widehat{Q}_{n,j}^{\Omega}(r)+\big[1-\widehat{Q}_{n,j}^{\Omega}(\mu)\big]\mathbb{I}\{R_{n}^{*}>r\}$ (16)
$=:\widehat{Q}_{n,j}^{\Omega}(r;\mu)$ (17)
@@
The crux of our hypothesis is that these conditional distributions from boosting can be aggregated across ancestors to approximate the climatological distribution $Q^{\Theta}(r,\mu)=P(R^{\ast}>r|R^{\ast}>\mu)$, where $\Theta$ is used to denote the ground truth that would be obtained from a long DNS. We specifically propose to aggregate the conditional CCDFs as a uniform mixture over ancestors, selecting one representative AST $A_{j_{n}}$ from each ancestor $n$ to best represent its alternate realities according to some selection rule (different rules will be evaluated thoroughly for the QG system in Sect. 6). We write the mixture as
$\widehat{Q}^{M}(r;\mu)=\frac{1}{N_{\text{short}}}\sum_{n=1}^{N_{\text{short}}}\widehat{Q}_{n,j_{n}}^{\Omega}(r;\mu),$ (18)
and call it the “MoCTail” estimator of $Q^{\Theta}(r,\mu)$, for “Mixture of Conditional Tails.”
@@
The recent works Noyelle (2024) and Bloin-Wibe et al. (2025) formulate a different estimator, which makes for an interesting comparison. Rather than summing $N_{\text{short}}$ tail CCDFs, each approximating a ratio of the form (6), they construct a single ratio by summing $N_{\text{short}}$ numerators and $N_{\text{short}}$ denominators.Translated into our own notation, this becomes
$\widehat{Q}^{P}(r;\mu)=\frac{\sum_{n=1}^{N_{\text{short}}}\widehat{Q}_{n,j_{n}}^{\Omega}(r)}{\sum_{n=1}^{N_{\text{short}}}\widehat{Q}_{n,j_{n}}^{\Omega}(\mu)}.$ (19)
We call this the “PoPTail” estimator of $Q^{\Theta}(r,\mu)$, for “Pool of Perturbed Tails.” Bloin-Wibe et al. (2025) do not model $R^{*}(\omega)$ parametrically, but instead use a standard Monte Carlo estimate $\widehat{Q}_{n,j}^{\Omega}(r)=$ (fraction of descendants exceeding $r$), which is probably necessary for their high-dimensional perturbations. However, we can convert the PoPTail estimator to our parametric version just by thinking in terms of CCDFs, hence the formulation in Eq. (19). The more important difference is that PoPTail avoids the potential degeneracy $\widehat{Q}^{\Omega}(\mu)=0$ by “pooling” non-extreme descendants together with extreme ones in the denominator.
One could argue for either estimator based on the validity of its underlying assumptions which are challenging to rigorously verify. Here we adopt a more openly empirical perspective in testing the skill of both.
An important advantage of both estimators is extensibility with respect to the dataset: if the variance is too high, one can always either generate new ancestors by extending the short DNS, or extend the range of ASTs sampled, or enlarge the ensemble at any ASTs deemed promising, without discarding the laborious samples already generated. This is unfortunately not the case with an algorithm like AMS, TEAMS, GKTL, or QDMC: because of the random rules by which ancestors are selected and new members generated, a completed run cannot be enlarged while retaining its estimation properties unless we are willing to do an entirely new additional run and combine estimates from multiple runs as was done in Ragone et al. (2018); Webber et al. (2019); Finkel and O’Gorman (2024). This results in waste during the fine-tuning process of calibrating TEAMS. For example, one might decide in retrospect that a TEAMS run was too aggressive in killing non-extreme simulations and raising the threshold and we can’t easily extend the run with a new set of hyperparameters. With boosting, we can simply go back, perturb those less-extreme simulations, and incorporate them into the dataset, without needing to re-generate everything. To make boosting competitive at sampling the highest levels of severity, we suspect it will be necessary to augment our current scheme with an iterative level-raising schedule, like TEAMS, but with less restriction on the sampling procedure.
2.3 Evaluating performance
We evaluate the MoCTail and PoPTail estimators $\widehat{Q}^{M}$ and $\widehat{Q}^{P}$ by comparing to the ground truth $Q^{\Theta}$ as estimated from a long DNS. DNS is in fact a trivial special case of ensemble boosting with $M=0$ (no descendants), reducing each summand of Eq. (18) and the numerator of Eq. (19) to $\mathbb{I}\{R_{n}^{*}>r\}$ and the denominator of Eq. (19) to $N_{\text{short}}$. Both estimators reduce to the same vanilla empirical CCDF in this case, and this is what we use to estimate $Q^{\Theta}$.
[…]
$\chi^{2}(\Delta\widehat{Q}\lVert\Delta Q^{\Theta})=\sum_{k=0}^{K-1}\frac{(\Delta Q^{\Theta}_{k}-\Delta\widehat{Q}_{k})^{2}}{\Delta Q^{\Theta}_{k}}$ (20)